# Prominent bump in the two-neutron separation energies of neutron-rich lanthanum isotopes revealed by high-precision mass spectrometry

### v1

@@ Abstract
We report on high-precision atomic mass measurements of ${}^{148\text{-}153}$a and 151Ce performed with the JYFLTRAP double Penning trap using the Phase-Imaging Ion-Cyclotron-Resonance technique. The masses of 152,153La were experimentally determined for the first time. We confirm the sharp kink in the two-neutron separation energies at the neutron number ${N=93}$ in the cerium (${Z=58}$) isotopic chain. Our precision mass measurements of the most exotic neutron-rich lanthanum (${Z=57}$) isotopes reveal a sudden increase in two-neutron separation energies from ${N=92}$ to ${N=93}$. Unlike in the cerium isotopic chain, the kink is not sharp but extends to ${N=94}$ forming a prominent bump. The gain in energy is about 0.4 MeV, making it one of the strongest changes in two-neutron separation energies over the whole chart of nuclides, away from nuclear shell closures. The results call for further studies to elucidate the structure of neutron-rich lanthanum isotopes.
@@
The masses of neutron-rich isotopes ranging from the barium ($Z=56$) to holmium ($Z=67$) isotopic chains have been extensively studied with the JYFLTRAP Penning trap [30, 31] and the Canadian Penning trap (CPT) [32, 33, 34, 35]. In the case of the neutron-rich barium isotopes, a relatively flat trend in the $S_{2n}$ values is observed. While the lanthanum (${Z=57}$) and cerium (${Z=58}$) isotopic chains exhibit a small kink in the $S_{2n}$ values at $N=93$, the lack of precise experimental data hinders drawing clear conclusions on possible structural effects. The masses of 149,151La are known with a limited precision of 200 and 440 keV [36, 37, 27], respectively, and the mass of 150La is known from a high-precision mass measurement performed at CPT [34]. In the case of 151Ce93 isotope, the deviation from a smooth trend could be explained by a presence of a long-lived isomer proposed in the ENSDF evaluation [38] but not observed in Ref. [32].
@@
The measurement was performed at the Ion Guide Isotope Separator On-Line (IGISOL) facility [65, 66]. The neutron-rich rare-earth isotopes were produced using $25$-MeV protons impinging on a $15$-mg/cm2-thick ${}^{\text{nat}}$ target. The fission fragments, mostly singly charged, were stopped in a helium-filled gas cell (at about 300 mbar) before extracted out through a sextupole ion guide [67]. The ions were further accelerated to $30q$ kV and separated based on their mass-to-charge ratio $m/q$ using a $55$ degree dipole magnet. The continuous beam was cooled and bunched with the Radiofrequency Quadrupole Cooler-Buncher [68] and sent to JYFLTRAP.
@@
Figure 2 shows two-neutron separation energies for the isotopic chains from caesium (${Z=55}$) to promethium (${Z=61}$) based on AME20 [27], including our new results on 148-153La and 151Ce, as well as the recent CPT measurements for cerium, praseodymium and promethium isotopic chains [35]. The values were obtained from the mass-excesses as $S_{2n}(Z,N)=\text{ME}(Z,N-2)-\text{ME}(Z,N)+2\text{ME}_{n}$, where $\text{ME}_{n}$ is the neutron mass excess. While the heavier isotopic chains in the rare-earth region exhibit prominent kinks at $N\approx 90$ due to onset of strong quadrupole deformation, the effect seems to vanish when moving to lower proton numbers according to AME20 values. This is in accordance with the energies of the first $2^{+}$ states, which decrease steeply at $N\geq 88$ for the even-even nuclei in the region but much less so for the lower-$Z$ isotopic chains of cerium and barium, which reach the minimum at larger $N$ [39].
@@
In this Letter we have reported on the high-precision mass measurements of 148-153La and 151Ce performed using the PI-ICR technique with the JYFLTRAP double Penning trap. With the new mass data, we could pin down the trends in the two-neutron separation energies $S_{2n}$. Our value for 151Ce confirms the small peak at ${N=93}$ in the cerium chain, coinciding with the transitional region from octupolar to quadrupolar deformation previously studied via $\gamma$-ray spectroscopy. For the lanthanum isotopes, our new values reveal a bump about 0.4-MeV in the $S_{2n}$ values at neutron numbers ${N=92-94}$, a unique feature within the neighboring isotopic chains which none of the studied mass models could reproduce. The change in the $S_{2n}$ trend agrees with the predicted onset of quadrupole deformation at ${N\geq 92}$ that could be potentially caused by the ${9/2^{+}[404]}$ proton extruder orbital. Such a proton extruder could be a plausible explanation for the localisation of the bump within the lanthanum and cerium isotopic chains, as for the other well-known kinks in two-neutron separation energies that span over many isotopic chains, neutron extruder orbitals play the main role. More extensive studies on exotic neutron-rich lanthanum isotopes at ${N\geq 90}$ would be highly welcome to elucidate the evolution of nuclear structure in the region.

### v2

@@ Abstract
We report on high-precision atomic mass measurements of ${}^{148\text{-}153}$a and 151Ce performed with the JYFLTRAP double Penning trap using the Phase-Imaging Ion-Cyclotron-Resonance technique. The masses of 152,153La were experimentally determined for the first time. We confirm the sharp kink in the two-neutron separation energies at the neutron number ${N=93}$ in the cerium (${Z=58}$) isotopic chain. Our precision mass measurements of the most exotic neutron-rich lanthanum (${Z=57}$) isotopes reveal a unexpected sudden increase in two-neutron separation energies from ${N=92}$ to ${N=93}$. Unlike in the cerium isotopic chain, the kink is not sharp but extends to ${N=94}$ forming a prominent bump. The gain in energy is about 0.4 MeV, making it one of the strongest changes in two-neutron separation energies over the whole chart of nuclides, away from nuclear shell closures. The results, correlated with a predicted onset of quadrupole deformation for $N\geq 92$, call for further studies to elucidate the structure of neutron-rich lanthanum isotopes.
@@
The masses of neutron-rich isotopes ranging from the barium ($Z=56$) to holmium ($Z=67$) isotopic chains have been extensively studied with the JYFLTRAP Penning trap [30, 31] and the Canadian Penning trap (CPT) [32, 33, 34, 35]. In the case of the neutron-rich barium isotopes, a relatively flat trend in the $S_{2n}$ values is observed. While the lanthanum (${Z=57}$) and cerium (${Z=58}$) isotopic chains exhibit a small kink in the $S_{2n}$ values at $N=93$, the lack of precise experimental data hinders drawing clear conclusions on possible structural effects. The masses of 149,151La are known with a limited precision of 200 and 440 keV [36, 37, 27], respectively, and the mass of 150La is known from a high-precision mass measurement performed at CPT [34]. In the case of 151Ce93 isotope, the deviation from a smooth trend could be explained by the presence of a long-lived isomer proposed in the ENSDF evaluation [38] but not observed in Ref. [32].
@@
The measurements were performed at the Ion Guide Isotope Separator On-Line (IGISOL) facility [65]. The neutron-rich rare-earth isotopes were produced using a $25$-MeV proton beam with up to 20 $\mu$A of intensity impinging on a $15$-mg/cm2-thick ${}^{\text{nat}}$ target. The fission fragments, mostly singly charged, were stopped in a helium-filled gas cell [66, 67] (at about 300 mbar) before being extracted out through a sextupole ion guide [68]. The ions were further accelerated to $30q$ kV and separated based on their mass-to-charge ratio $m/q$ using a $55$ degree dipole magnet. The continuous beam was cooled and bunched with the Radiofrequency Quadrupole Cooler-Buncher [69] and sent to JYFLTRAP.
@@
Figure 2 shows two-neutron separation energies for the isotopic chains from caesium (${Z=55}$) to promethium (${Z=61}$) based on AME20 [27], including our new results on 148-153La and 151Ce, as well as the recent CPT measurements for cerium, praseodymium and promethium isotopic chains [35]. The values were obtained from the mass-excesses as $S_{2n}(Z,N)=\text{ME}(Z,N-2)-\text{ME}(Z,N)+2\text{ME}_{n}$, where $\text{ME}_{n}=8071.3181(4)$ keV is the neutron mass excess [27]. While the heavier isotopic chains in the rare-earth region exhibit prominent kinks at $N\approx 90$ due to onset of strong quadrupole deformation, the effect seems to vanish when moving to lower proton numbers according to AME20 values. This is in accordance with the energies of the first $2^{+}$ states, which decrease steeply at $N\geq 88$ for the even-even nuclei in the region but much less so for the lower-$Z$ isotopic chains of cerium and barium, which reach the minimum at larger $N$ [39].
@@
In this Letter we have reported on the high-precision mass measurements of 148-153La and 151Ce performed using the PI-ICR technique with the JYFLTRAP double Penning trap. With the new mass data, we could pin down the trends in the two-neutron separation energies $S_{2n}$. Our value for 151Ce confirms the small peak at ${N=93}$ in the cerium chain, coinciding with the transitional region from octupolar to quadrupolar deformation previously studied via $\gamma$-ray spectroscopy. Our measurements extend the limits of the known lanthanum masses and reveal around 0.4-MeV bump in the $S_{2n}$ values at neutron numbers ${N=92-94}$, a unique feature within the neighboring isotopic chains not reproduced by any studied mass models. The observed change in the $S_{2n}$ trend agrees with the predicted onset of quadrupole deformation at ${N\geq 92}$ that could be potentially caused by the ${9/2^{+}[404]}$ proton extruder orbital. This could be a plausible explanation for the localisation of the bump, as analogously neutron extruder orbitals play the main role for the other well-known kinks in two-neutron separation energies. Finally, we note that the mass evolution of neutron-rich isotopes further away from stability is crucial for astrophysical calculations, as being one of the key inputs in the determination of neutron-capture reaction rates. Specifically, the isotopes reported in this work and the possible structural changes in their vicinity were predicted to be of pivotal importance in the $r$-process [95, 96, 97, 98]. A better understanding of the nuclear structure in this region is thus of broader interest for the physics community and calls for more extensive experimental and theoretical studies.